# Current density distribution for the quantum Hall effect

### v2

@@ Abstract
Our microscopic understanding of the integer quantum Hall effect is still incomplete. For decades, there has been a controversial discussion about “where the current flows” if the Hall resistance is quantized. Here, we qualitatively analyze the current density distribution in a Hall bar based on the screening properties of a two-dimensional electron system in the quantum Hall regime. Beyond previous publications, we include a closed loop persistent current that exists inside a Hall bar if the Hall resistance is quantized. We find, that the persistent current density decreases with increasing Hall voltage. Accounting for this dependence, we find, that the current flows in the opposite directions along opposite edges of the Hall bar, while the imposed current flows unidirectionally and only on the side of the Hall bar connected with its higher electrical potential edge.
@@ I Introduction
The integer quantum Hall effect (QHE) is one of the most fundamental and important phenomena in condensed matter physics. The quantized Hall resistance serves as a reference for the SI (Système International) unit system [1, 2], while new applications based on the QHE are likely to emerge in the context of quantum technologies. To realize such applications, it is important to understand the microscopic distributions of the electrostatic potential as well as the current density. For instance, the electrostatic details influence the stability of the quantized Hall plateaus, while the current density distribution also affects the quantum mechanical phase of the charge carriers [3]. For a known confinement potential of a hypothetical Hall bar without disorder, the distribution of the electrostatic potential can be numerically predicted with a high accuracy [4]. Nevertheless, the question of how to infer from the electrostatics to the microscopic distribution of the current in the case of the QHE is still the subject of controversal debate. Clarity will not only pave the way for new applications but also enhance our understanding of the fractional, spin or anomalous quantum Hall effects.
The text book explanation of the QHE, namely the Landauer-Büttiker picture, suggests that for the plateaus of the QHE all current flows inside one-dimensional channels along the edges of the Hall bar [5, 6]. This view is challanged by experiments [7, 8, 9, 10] and theoretical calculations taking into account the Coulomb screening of the charge carriers [11, 12, 4]. Both show, that for specific regions of the plateaus the Hall voltage drops continuously across the Hall bar. Since current is generated by electric fields, this indicates the existence of bulk currents flowing far away from the edges of the Hall bar. Recently, such bulk currents have been directly measured in a narrow, high mobility Hall bar [13]. Recent multiterminal experiments probing the distribution of $I_{\text{impo}}$ into various grounded contacts, however, suggest that for the entire plateaus $I_{\text{impo}}$ follows a chiral trajectory, i.e., along the sample edge from contact to contact, while in between plateaus, $I_{\text{impo}}$ is unidirectional as described by the Drude model [14].
In the present article, we present a qualitative model for the current density distribution, which explains all these findings. We restrict our analysis to the plateaus of the QHE within the scope of the screening theory of the QHE [15, 16, 17, 18, 11, 12, 4]. Our analysis builds on the calculations of the screening theory for the electrostatic potential distribution similar as in a previous prediction of the current density distribution [19], but we include a closed loop persistent current, $I_{\text{pers}}$, that was completely neglected in Ref. [19]. The persistent current is confined inside the Hall bar and exists in addition to the current, $I_{\text{impo}}$, imposed on the Hall bar through ohmic contacts, if the Hall resistance is quantized. $I_{\text{pers}}$ directly depends on $I_{\text{impo}}$, which affects the distribution of the overall current, $I_{\text{impo}}+I_{\text{pers}}$, inside the Hall bar. Taking this dependence into account, we find, that the current flow is chiral, in the sense that the current flows in opposite directions along the opposite edges of a Hall bar. Subtracting the persistent current density from the overall current yield the imposed current. We find, that the imposed current flows unidirectionally and is concentrated on the side of the Hall bar with the higher electrical potential. It vanishes along the edge with the lower electrical potential. This result contradicts Ref. [19], which predicted that the imposed current flows unidirectionally on both sides of the Hall bar.
Non-local nature of the QHE and persistent current
@@
The ballistic trajectory of a charged particle exposed to a magnetic field is chiral, a fundamental property related with the axial symmetry of the magnetic field vector. In solids, however, boundary conditions and momentum scattering can alter the impact of the axial symmetry and obscure the chirality of trajectories. This is clearly the case for the classical Hall effect described by the Drude model, which predicts a homogeneous unidirectional current density. The Drude model uses a relaxation time ansatz, which is local in nature and justified in the diffusive regime for $\mu B\lesssim 1$, where $\mu$ is the mobility [14].
The situation is different in the regime of the QHE with $\mu B\gg 1$ (and low enough temperature), where the current density is governed by the Landau quantization [20]. In the two-dimensional electron system (2DES) of a Hall bar and for the plateaus of a quantized Hall resistance, it gives rise to regions in which a local energy gap suppresses carrier scattering and Coulomb screening. In such regions, the sloped electrostatic potential caused by the lateral confinement of the Hall bar, provides the driving force of a dissipationless persistent current. It flows along a chiral loop along the sample edge even inside an unbiased Hall bar [16]. Applying a voltage between ohmic contacts of the Hall bar, breaks the equilibrium by imposing an external current, $I_{\text{impo}}$. The latter generates the Hall voltage, $V_{\text{H}}=R_{\text{H}}I_{\text{impo}}$, dropping between the sample edges, where $R_{\text{H}}=h/(e^{2}\nu)$ is the Hall resistance with filling factor $\nu$, elementary charge $e$ and Planck constant $h$. Throughout quantized plateaus of the integer QHE, the filling factor remains integer within the extended regions of suppressed Coulomb screening and, $I_{\text{impo}}$, too, flows dissipationless through the Hall bar. The dissipation implied by the Hall resistance happens within “hot-spots” near the current carrying contacts [21, 22]. Both, “hot-spots” and persistent current are signatures of a non-local nature of the QHE. The current density distribution is strongly affected by this non-locality including a cross dependence $I_{\text{pers}}(I_{\text{impo}})$. This means in particular, that in order to determine the distribution of $I_{\text{impo}}$, it is necessary to consider the complete current density, $j=j_{\text{impo}}+j_{\text{pers}}$.
@@ II Screening Theory
Exposing a 2DES to a high perpendicular magnetic field leads to the collaps of the density of states (DOS) into discrete LLs, the physical cause of the QHE. The Landauer-Büttiker picture of the QHE [5, 6], neglects the Coulomb interaction between charge carriers. However, the energy gap between LLs would then lead to displacements of the electrons causing high local charges [15]. The screening theory fixes this problem by taking into account the direct Coulomb interaction between carriers [15, 16, 17, 18, 11, 12, 4]. This way, it accounts for the tendency of the carriers to minimize the free energy by redistribution (thereby avoiding high local charges) [15]. Nevertheless, in some regions of the 2DES, all LLs still end up to be either completely filled or empty, such that a local energy gap,
$\Delta E=\left\{\begin{array}[]{cl}g\mu_{\text{B}}B&\text{; for odd }\nu\\ \hbar\omega_{\text{c}}-g\mu_{\text{B}}B&\text{; for even }\nu\end{array}\right$.\,, (1)
remains, where $g$ is the Landee g-factor, $\mu_{\text{B}}$ the Bohr magneton, $\nu$ the filling factor and $\omega_{\text{c}}$ the cyclotron frequency. For $k_{\text{B}}T\ll\Delta E$ and low enough disorder, this gap locally suppresses scattering of electrons because of a lack of accessible unoccupied states. In a sense, the electrons behave like an incompressible liquid in these regions, which are therefore called incompressible strips (ICSs). The screening theory proposes a fragmentation of the Hall bar in compressible perfectly screened regions, separated by the unscreened ICSs [11, 12, 4]. Inside a compressible region, one partly filled LL is pinned to the locally constant chemical potential $\mu(y)$, hence the energies of all LLs, $\varepsilon_{\nu}$, are constant. Consequently, inside the unscreened ICSs, $\mu(y)$ and $\varepsilon_{\nu}(y)$ vary by the amount of the energy gap.
In Fig. 1,
@@
we sketch a typical situation as it is predicted by the screening theory near the low magnetic field end of the first plateau of the QHE, i.e., for $\nu=1$ inside the ICSs (assuming a symmetric confinement potential). In Fig. 1(a) we depict the equilibrium situation for $I_{\text{impo}}=0$ and $\mu_{0}=const$, while in Fig. 1(b) we show a linear approximation for $I_{\text{impo}}\neq 0$ and varying $\mu_{0}(y)$. The energies, $\mu_{0}(y)$ and $\varepsilon(y)\equiv\varepsilon_{1}(y)$ vary only within the ICSs. However, their actual values depend on $I_{\text{impo}}$ even inside the compressible region between the ICSs.
The early formulations of the screening theory were content with analytical and first self-consistent approximations [15] of the distribution of the electron density based on the Poisson equation and using the Thomas-Fermi approximation [16]. Later, the idea was refined by applying a self-consistent non-linear Hartree-type calculation and accounting for the quantum mechanical wave properties of the electrons by means of a Gaussian broadening of the LLs [11, 12]. Recently, (for high magnetic fields) the electrostatic results where confirmed for low integer filling factors within a full self-consistent approach of the quantum-electrostatic problem without using the Thomas-Fermi approximation [23].
Electrostatics of the QHE
@@
We consider a Hall bar with edges at $y=\pm b$, defined for the carrier density $n_{\text{s}}>0$ for b<y<b-b<y<b, see Fig. 1. A current, $I_{\text{impo}}$, can be imposed between source (S) and drain (D) contacts, ideally positioned far away (at $x=\mp\infty$) along the vertical axis in the sketches in Fig. 1. In our sketches, we show a typical case as it occurs for a quantized Hall resistance near the low magnetic field end of the plateau. The local filling factor inside the ICSs is $\nu(y)=1$. If the current flows inside the ICSs, it is carried by a single (spin polarized) LL with energy $\varepsilon(y)$. The energy of the second LL is higher by the Zeeman splitting $\Delta E=g\mu_{\text{B}}B$. As sketched in Fig. 1, there exist compressible strips between the edges and the ICSs, where the first LL is pinned to the chemical potential ($\nu<1$) and a compressible bulk region between the ICSs, where the second LL is pinned to the chemical potential and partly occupied ($1<\nu<2$). For simplicity, we neglect many body correlations, interference effects or details at the interface between compressible regions and the ICSs. Further, we consider $I_{\text{impo}}$ to be small compared to the break-down current [24], such that a quantized plateau and ICSs can exist, and assume an idealized 2DES at zero temperature without disorder and a symmetric confinement potential. Moderate disorder effects [25] as well as asymmetric confinement [26] would have quantitative consequences but will not affect the qualitative findings of this article.
Geometry of the ICSs – linear approximation
@@
In Fig. 1, we assumed that the width, $w_{\text{ICS}}$, of the ICSs is independent of $I_{\text{impo}}$. This corresponds to the linear approximation of the screening theory for a symmetric Hall bar. A more accurate examination reveals a non-linear relation between the imposed current and the induced change of the electrostatic potential, which can be accounted for in a self-consistent numerical calculation [4]. Qualitatively, the non-linear dependence is already present in a simple electrostatic approximation considering the local energy balance. It considers the Coulomb energy, $E_{\text{C}}$, caused by charge separation across an ICS because of its constant carrier density. Requiring $E_{\text{C}}=\Delta E$, it predicts for the equilibrium $w_{\text{ICS}}\propto\sqrt{\Delta E}$ [15]. Applying the same argument for $I_{\text{impo}}\neq 0$ results (in the limit of individual edge ICSs) in $w_{\text{ICS}}\propto\sqrt{\Delta E+\Delta\mu}$, where $\Delta\mu$ is the change of $\mu(y)$ across an ICS. Because of the confinement, $\Delta E$ is negative on one and positive on the other side of the Hall bar, cf. Fig. 1(a), while the change of the chemical potential across an ICS, $\Delta\mu$, has the identical sign for all ICSs, see Fig. 1(b). Therefore, $w_{\text{ICS}}$ is reduced on one and increased on the other side of the Hall bar. To keep our discussion simple, in Fig. 1 we stick to the linear approximation, which predicts constant $w_{\text{ICS}}$ and corresponds to the first loop of the self-consistent iterations [4]. However, the actual increase of $\Delta E\propto B$ as $B$ is increased along a plateau of the QHE leads to the formation of a bulk ICS towards the high magnetic field end of a plateau [12]. Since this bulk ICS explains the measured bulk current [13] and bulk variations of $\mu(y)$ on the plateaus [10], we consider this limit in the right hand column of Fig. 2. Note that, considering the non-linear screening behavior quantitatively would require self-consistent numerical calculations as in Ref. [12] but wouldn’t change our qualitative arguments below.

### v3

@@ Abstract
Our microscopic understanding of the integer quantum Hall effect is still incomplete. For decades, there has been a controversial discussion about “where the current flows” if the Hall resistance is quantized. Here, we analyze the current density distribution in a Hall bar based on the screening properties of a two-dimensional electron system in the quantum Hall regime. Compared to previous publications, we consider the non-local nature of the quantum Hall effect and include a related persistent current that exists inside a Hall bar if the Hall resistance is quantized. We find, that the persistent current density decreases with increasing Hall voltage. Accounting for this dependence, we find, that the current flows in the opposite directions along opposite edges of the Hall bar, while the share of the current imposed through the contacts flows along the edge of the Hall bar with the higher chemical potential.
@@ I Introduction
The integer quantum Hall effect (QHE) is one of the most fundamental and important phenomena in condensed matter physics. The quantized Hall resistance serves as a reference for the SI (Système International) unit system BIPM (2019); Delahaye and Jeckelmann (2003), while new applications based on the QHE are likely to emerge in the context of quantum technologies. To realize such applications, it is important to understand the microscopic distributions of the electrostatic potential as well as the current density. For instance, the electrostatic details influence the stability of the quantized Hall plateaus, while the current density distribution also affects the quantum mechanical phase of the charge carriers Ofek et al. (2010). For a known confinement potential of a hypothetical Hall bar without disorder, the distribution of the electrostatic potential can be numerically predicted with a high accuracy Gerhardts (2008). Nevertheless, the question of how to infer from the electrostatics to the microscopic distribution of the current in the case of the QHE is still the subject of controversal debate. Clarity will not only pave the way for new applications but also enhance our understanding of the fractional, spin or anomalous quantum Hall effects.
The text book explanation of the QHE, namely the Landauer-Büttiker picture, suggests that for the plateaus of the QHE all current flows inside one-dimensional channels along the edges of the Hall bar Büttiker (1986, 1988). This view is challanged by experiments Fontein et al. (1991); Yacoby et al. (1999); McCormick et al. (1999); Weitz et al. (2000) and theoretical calculations taking into account the Coulomb screening of the charge carriers Güven and Gerhardts (2003); Siddiki and Gerhardts (2003, 2004); Gerhardts (2008). Scanning probe experiments Weitz et al. (2000) and the so-called screening theory Gerhardts (2008) suggest that near the low magnetic field end of a plateau the entire Hall potential drops within narrow strips near the edges of the sample. However, as the magnetic field is increased these strips of finite potential gradient widen and extend into the Hall bar, until —near the high magnetic field end of the plateau— the Hall potential drops continuously across the entire Hall bar. If the local current density is linked to the gradient of the potential, the scanning probe experiments indicate a continues transition from edge to bulk currents as the magnetic field is increased for each plateau of the QHE Güven and Gerhardts (2003).
A direct non-invasive measurement of the current density distribution in the regime of the QHE is impossible. Nevertheless, carefully designed transport experiments allow one to study important aspects of the current flow. For example, by using a miniature ohmic contact in the center of a high mobility Hall bar as a current probe, recently the predicted transition from edge to bulk current was observed while the magnetic field was increased along the plateaus Sirt et al. (2025a). In another experiment Sirt et al. (2025b), multiterminal measurements were used to test whether current flows unidirectionally across the entire Hall bar or follows a chiral trajectory along the sample edge. The measurements suggest that for the entire plateaus the current follows a chiral trajectory, while in between plateaus, the current flows unidirectional as described by the Drude model Sirt et al. (2025b).
In summary, the electrostatic potential distribution predicted by the screening theory confirms for the plateaus of the QHE the results of scanning probe measurements Weitz et al. (2000). Transport experiments corroborate that for the plateaus of the QHE the current flows where the potential gradient is finite. This gives rise to chiral current flow, including both, edge and bulk currents depending on the magnetic field Sirt et al. (2025a, b).
In the present article, we present a simple model for the current density distribution, which consistently explains these experimental findings. For simplicity, we consider a two-terminal sample consisting of a two-dimensional electron system (2DES) connected at both ends to ohmic contacts. A current $I_{\text{impo}}$ is imposed via these contacts and flows through the 2DES. Poor ohmic contacts or additional electrically floating contacts, typically used as voltage probes, can influence the stability of the QHE but do not change the general physics discussed here. We avoid such implications, by restricting our analysis to the plateau regions of the QHE.
Non-local nature of the QHE
@@
The ballistic trajectory of a charged particle exposed to a magnetic field is chiral, a fundamental property related with the axial symmetry of the magnetic field vector. In solids, however, boundary conditions and momentum scattering can alter the impact of the axial symmetry and obscure the chirality of the trajectories of carriers. This is clearly the case for the classical Hall effect described by the Drude model, as it predicts a homogeneous unidirectional current density. The Drude model uses a relaxation time ansatz, which is local in nature and well justified in the diffusive regime for $\mu B\lesssim 1$ Sirt et al. (2025b), where $\mu$ is the mobility and $B$ the magnitude of the magnetic field component perpendicular to the plane of the 2DES.
The situation is different in the regime of the QHE with $\mu B\gg 1$ (and $k_{\text{B}}T\ll\hbar\omega_{\text{c}}$ with the Boltzman constant $k_{\text{B}}$, $\hbar=h/2\pi$, the Planck constant $h$ and the cyclotron frequency $\omega_{\text{c}}$), where the current density is governed by the Landau quantization von Klitzing et al. (1980). In the 2DES of a Hall bar, for the plateaus of a quantized Hall resistance a local energy gap gives rise to regions of suppressed carrier scattering and Coulomb screening, such that within these regions current can flow virtually without dissipation Chklovskii et al. (1993).
An externally applied current $I_{\text{impo}}$ generates the Hall voltage, $V_{\text{H}}=R_{\text{H}}I_{\text{impo}}$, dropping between the sample edges, where $R_{\text{H}}=h/(e^{2}\nu)$ is the Hall resistance and $\nu$ the local filling fraction of the Landau levels (LLs). According to the screening theory, $\nu$ remains integer within the extended regions of suppressed Coulomb screening, such that throughout the plateaus of the integer QHE $I_{\text{impo}}$ flows dissipationless through the entire Hall bar. The dissipation related with the Hall resistance happens within “hot-spots” near the current carrying contacts Klaß et al. (1991); Komiyama et al. (2006).
The “hot-spots” and the dissipationless current within the Hall bar are signatures of a non-local nature of the QHE.
Persistent current of the QHE
The non-local nature of the QHE has a surprising consequence, namely a macroscopic dissipationless current density that persists even in equilibrium, i.e., for $I_{\text{impo}}=0$ Chklovskii et al. (1993); Wexler and Thouless (1994); Geller and Vignale (1994, 1995); Komiyama and Hirai (1996); Güven and Gerhardts (2003). While this additional current is often labeled “equilibrium current”, here we call it persistent current, $I_{\text{pers}}$. Note, that this macroscopic $I_{\text{pers}}$ is not related with the persistent current in mesoscopic rings Gefen et al. (1984). Instead, it is driven by the effective electric fields corresponding to the gradient of the confinement potential near the edges of a Hall bar. Such local electric fields exist due to the above mentioned energy gap between LLs, which effectively prevents screening.
A previous calculation of $I_{\text{impo}}$ in Rev. Güven and Gerhardts, 2003 applies a local version of Ohm’s law to the distribution of the chemical potential predicted by the screening theory Chklovskii et al. (1992, 1993); Fogler and Shklovskii (1994); Lier and Gerhardts (1994); Siddiki and Gerhardts (2003, 2004); Gerhardts (2008) and thereby completely neglects $I_{\text{pers}}$. It predicts that $I_{\text{impo}}$ generally flows in the same direction along both edges of the Hall bar. However, in this article, we show, that $I_{\text{pers}}$ depends on $I_{\text{impo}}$, such that for the case of a finite $I_{\text{impo}}$, the neglect of $I_{\text{pers}}$ does not yield the correct current density distribution.
Our analysis presented in this article builds on the predictions of the screening theory, in common with Ref. Güven and Gerhardts, 2003. However, we appretiate the non-local nature of the QHE and explicitly include $I_{\text{pers}}$ in our model. Driven by the gradient of the confinement potential near the sample edges, $I_{\text{pers}}$ flows in opposite directions along the opposing egdes of a Hall bar. Thereby, $I_{\text{pers}}$ is carried by electrons at energies below the chemical potentials of both leads. Clearly, it cannot contribute to the imposed current, $I_{\text{impo}}$ (carried by electrons at energies between the chemical potentials of both contacts). However, we find that $I_{\text{pers}}$ decreases as $I_{\text{impo}}$ is increased. Accounting for this dependence, $I_{\text{pers}}(I_{\text{impo}})$, we determine the distributions of $I_{\text{pers}}$ and $I_{\text{impo}}$ inside the Hall bar.
We find, that the overall current flow is chiral, in the sense that the current flows in opposite directions along the opposite edges of a Hall bar, while the imposed current is concentrated on the side of the Hall bar with the higher electrical potential, i.e., $I_{\text{impo}}$ always vanishes along the edge with the lower potential. Our results are in agreement with recent experimental observations Sirt et al. (2025b), but contradict the claim of Ref. Güven and Gerhardts, 2003, that $I_{\text{impo}}$ would flow in the same direction along both edges of the Hall bar.
@@ II Screening Theory
Exposing a 2DES to a high perpendicular magnetic field leads to the collaps of the density of states (DOS) into discrete LLs, the physical cause of the QHE. The Landauer-Büttiker picture of the QHE Büttiker (1986, 1988), neglects the Coulomb interaction between charge carriers. However, the energy gap between LLs would then lead to displacements of the electrons causing high local charges Chklovskii et al. (1992). The screening theory fixes this problem by taking into account the direct Coulomb interaction between carriers Chklovskii et al. (1992, 1993); Fogler and Shklovskii (1994); Lier and Gerhardts (1994); Siddiki and Gerhardts (2003, 2004); Gerhardts (2008). This way, it accounts for the tendency of the carriers to minimize the free energy by redistribution (thereby avoiding high local charges) Chklovskii et al. (1992). Nevertheless, in some regions of the 2DES, all LLs still end up to be either completely filled or empty, such that a local energy gap,
$\Delta E_{\nu}=\left\{\begin{array}[]{cl}g\mu_{\text{B}}B&\text{; for odd }\nu\\ \hbar\omega_{\text{c}}-g\mu_{\text{B}}B&\text{; for even }\nu\end{array}\right$.\,, (1)
remains, where $g$ is the Landee g-factor and $\mu_{\text{B}}$ the Bohr magneton. For $k_{\text{B}}T\ll\Delta E_{\nu}$ and low enough disorder, this gap locally suppresses scattering of electrons because of a lack of accessible unoccupied states. In a sense, the electrons behave like an incompressible liquid in these regions, which are therefore called incompressible strips (ICSs). These are the regions of suppressed scattering already mentioned in Sec. I. In more detail, the screening theory proposes a fragmentation of the Hall bar in compressible perfectly screened regions, separated by the unscreened ICSs Siddiki and Gerhardts (2003, 2004); Gerhardts (2008). Inside a compressible region, one partly filled LL is pinned to the locally constant chemical potential $\mu(y)$, hence the energies of all LLs, $\varepsilon_{\nu}$, are also constant. Consequently, inside the unscreened ICSs, both, $\mu(y)$ and $\varepsilon_{\nu}(y)$ vary, while the filling factor is constant.
The early formulations of the screening theory were content with analytical and first self-consistent approximations of the distribution of the electron density based on the Poisson equation Chklovskii et al. (1992) and using the Thomas-Fermi approximation Chklovskii et al. (1993). Later, the idea was refined by applying a self-consistent non-linear Hartree-type calculation and accounting for the quantum mechanical wave properties of the electrons by means of a Gaussian broadening of the LLs Siddiki and Gerhardts (2003, 2004). Recently, (for high magnetic fields) the electrostatic results where confirmed for low integer filling factors within a full self-consistent approach of the quantum-electrostatic problem without using the Thomas-Fermi approximation Armagnat and Waintal (2020).
Electrostatics of the QHE
@@
We consider a Hall bar with edges at $y=\pm b$, defined for the carrier density $n_{\text{s}}>0$ for b<y<b-b<y<b. In Fig. 1, (assuming a symmetric confinement potential) we sketch the first two LLs (solid black line for the first and dashed black line for the second LL) as they are predicted by the screening theory near the low magnetic field end of the first plateau of the QHE. The energy gap between the first two LLs is $\Delta E_{1}=g\mu_{\text{B}}B$. In the considered limit, two narrow ICSs with filling factor $\nu=1$ form near the edges of the Hall bar. The energy of the LLs varies within the ICSs only. A current, $I_{\text{impo}}$, can be imposed between source (S) and drain (D) contacts, ideally positioned far away (at $x=\mp\infty$ along the vertical axis in the sketches in Fig. 1). In Fig. 1(a) we depict the equilibrium situation for $I_{\text{impo}}=0$ and constant $\mu_{0}(y)$, while in Fig. 1(b) we show a linear approximation for $I_{\text{impo}}\neq 0$. For our example of two ICSs with $\nu=1$, the energy variation of the LLs across the two ICSs is $\Delta\varepsilon_{\pm}=\pm\Delta E_{1}+\Delta\mu_{\pm}$, where $\mu_{\pm}$ correspond to the individual variations of the Hall potential in the two ICSs for $y\gtrless 0$ with $\Delta\mu_{+}+\Delta\mu_{-}=-eV_{\text{H}}$. Within the compressible strips between the edges and the ICSs, the first LL is pinned to the chemical potential ($\nu<1$). Within the compressible bulk region between the ICSs, the second LL is pinned to the chemical potential and partly occupied ($1<\nu<2$). In Sec. III below, we show that the current flows inside the ICSs. In our example in Fig. 1, it is carried by a single (spin polarized) LL.
For simplicity, we neglect many body correlations, interference effects or details at the interface between compressible regions and the ICSs. Further, we consider $I_{\text{impo}}$ to be small compared to the break-down current Panos et al. (2014), such that a quantized plateau and ICSs can exist, and assume an idealized 2DES at zero temperature without disorder and a symmetric confinement potential. Moderate disorder Siddiki and Gerhardts (2007) as well as asymmetric confinement Siddiki et al. (2010) would lead to corrections of the current distribution and influence the extent of the plateaus as well as the breakdown of the QHE. However, such influences will not affect the main findings of this article.
Geometry of the ICSs – non-linear effects and linear approximation
@@
In Fig. 1, we assumed that the width, $w_{\text{ICS}}$, of the ICSs is independent of $I_{\text{impo}}$. This corresponds to the linear approximation of the screening theory for a symmetric Hall bar. A more accurate examination reveals a non-linear relation between the imposed current and the induced change of the electrostatic potential, which can be accounted for in a self-consistent numerical calculation Gerhardts (2008). The non-linear dynamics of the screening is already present in a simple electrostatic approximation considering the local energy balance. It considers the Coulomb energy, $E_{\text{C}}$, caused by charge separation across an ICS, which occurs for a constant filling factor inside of the ICSs. In steady state, the Coulomb energy equals the local energy variation of the LLs, $E_{\text{C}}=\Delta\varepsilon_{\pm}$, which allows us to predict the width of an ICS, $w_{\text{ICS}}$. As long as the gradient of the (unscreened) confinement potential is constant within an ICS, $E_{\text{C}}\propto w_{\text{ICS}}^{2}$ is a reasonable approximation and we find $w_{\pm}\propto\sqrt{|\Delta\varepsilon_{\pm}}|$ Chklovskii et al. (1992). For $I_{\text{impo}}\neq 0$ this result implies $w_{+}>w_{-}$, because $|\Delta\varepsilon_{+}|-|\Delta\varepsilon_{-}|=|eV_{\text{H}}|$, cf. Fig. 1. In effect, as we increase $|I_{\text{impo}}|$, the width $w_{\text{ICS}}$ of the ICS on the high potential side of the Hall bar is increased, while the width of the ICS on the low potential side of the Hall bar it is reduced. To keep our discussion simple, in Fig. 1 we stick to the linear approximation, which corresponds to the first loop of the self-consistent iterations and results in constant $w_{\text{ICS}}$ Gerhardts (2008).
Due to the curvature of the confinement potential, the geometry of the ICS also depends on the magnetic field: The carrier density and the filling factor increase with increasing distance from the sample edge. At the same time, at a given position inside the Hall bar, the filling factor decreases as $B$ is increased. Because, the integer filling factor has to remain constant inside the ICSs, as $B$ is increased each ICS will move from the edge towards the center of the Hall bar Chklovskii et al. (1992). At the same time, its width increases because the absolute value of gradient of the confinement potential becomes smaller towards the center of the Hall bar. As $B$ is increased along a plateau of the QHE, this dynamics yields a transition from narrow edge ICSs to a wide bulk ICS towards the high magnetic field end of a plateau Siddiki and Gerhardts (2004). In Fig. 2 we compare the extreme limits, narrow ICSs near the edges of the Hall bar in the left column versus a single bulk ICS in the right column. The bulk currents measured in Ref. Sirt et al., 2025a and the variations of $\mu(y)$ reported in Weitz et al. (2000) confirm the predicted Siddiki and Gerhardts (2004) formation of a bulk ICS. Note that, considering the non-linear screening behavior quantitatively would require self-consistent numerical calculations as done in Ref. Siddiki and Gerhardts, 2004. While this would modify the detailed current distribution, the prediction of the following calculation remains otherwise valid.